# The Multi-Query Paradox in Zeroth-Order Optimization

### v1

@@ 6 Conclusions
This work addresses two core questions in zeroth-order optimization: how to aggregate multi-query information and how to allocate a query budget. We analyze two strategies: simple averaging (ZO-Avg) and a principled Projection Alignment (ZO-Align) method derived from local surrogate minimization. Our analysis, validated across numerous theoretical settings and experiments, reveals a stark dichotomy. We prove the optimal strategy for ZO-Avg is always a sequential random search, while ZO-Align’s performance consistently improves as more queries are used, culminating in a full-subspace estimation. This resolves the multi-query paradox, showing the choice is not about an intermediate query size, but about committing to an aggregation method that determines one of these two canonical algorithms.
Appendix A Appendix
A.1 Proofs
Proof of Proposition 3.1
@@ A.1 Proofs
Proof of Proposition 3.2
@@
Proof of Theorem 4.1
@@
Using the properties $\mathbb{E}[\hat{g}_{AVG}(x_{t})]=\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{AVG}(x_{t})\|^{2}]=\frac{q_{t}+d+1}{q_{t}}\|\nabla f(x_{t})\|^{2}$, and setting the optimal step size $\eta_{t}^{*}=\frac{q_{t}}{L(q_{t}+d+1)}$, we obtain the one-step progress:
@@
By $\mu$-strong convexity, $\|\nabla f(x_{t})\|^{2}\geq 2\mu(f(x_{t})-f(x_{*}))$. Let $\delta_{t}=E[f(x_{t})-f(x_{*})]$. Taking the full expectation and applying the convexity condition gives:
$\delta_{t+1}\leq\left(1-\frac{\mu q_{t}}{L(q_{t}+d+1)}\right)\delta_{t}$
@@
$\delta_{T}\leq\left(1-\frac{\mu q_{T-1}}{L(q_{T-1}+d+1)}\right)\delta_{T-1}$
$\leq\left(1-\frac{\mu q_{T-1}}{L(q_{T-1}+d+1)}\right)\left(1-\frac{\mu q_{T-2}}{L(q_{T-2}+d+1)}\right)\delta_{T-2}$
$\dots\leq\left(\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{L(q_{t}+d+1)}\right)\right)\delta_{0}$
@@
For the second part, we must find the sequence $\{q_{t}\}_{t=0}^{T-1}$ that minimizes the final error bound under the budget constraint. This is equivalent to minimizing the total contraction factor. Let $\rho_{t}=\frac{\mu q_{t}}{L(q_{t}+d+1)}$. We solve the following optimization problem:
@@
$\max_{\{q_{t}\},T}\sum_{t=0}^{T-1}\frac{\mu}{L}h(q_{t})\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
@@
As in A.1, we establish the one-step contraction factor. Using the properties $\mathbb{E}[\hat{g}_{ALG}(x_{t})]=\frac{q_{t}}{d}\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{ALG}(x_{t})\|^{2}]=\frac{q_{t}}{d}\|\nabla f(x_{t})\|^{2}$ with the optimal step size $\eta_{t}^{*}=1/L$, we find:
@@
Applying the $\mu$-strong convexity condition and taking the full expectation yields:
$E[f(x_{t+1})-f(x_{*})]\leq\left(1-\frac{\mu q_{t}}{Ld}\right)(E[f(x_{t})]-f(x_{*}))$
@@
$\min_{\{q_{t}\},T}\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{Ld}\right)\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
Let $c=\mu/(Ld)$. The objective function $F(q_{0},\dots,q_{T-1})=\prod_{t=0}^{T-1}(1-cq_{t})$ is a product of linear terms. This function is concave over the feasible region (the simplex defined by the budget constraint). A fundamental result in convex optimization is that the minimum of a concave function over a compact convex set is achieved at an extreme point (a vertex) of the set.

### v3

@@ 7 Conclusions
This work addresses two core questions in zeroth-order optimization: how to aggregate information from multiple queries and how to allocate a fixed query budget. By analyzing simple averaging ZO-Avg and the projection-based ZO-Align estimator across strongly convex, convex, non-convex, and stochastic settings, we show that aggregation and allocation are tightly coupled and lead to a sharp dichotomy. Under a fixed budget, the optimal strategy for ZO-Avg is always to use single-query steps, whereas ZO-Align can benefit from larger query blocks. Experiments validate the theoretical analysis. Our results resolve the multi-query paradox by clarifying that the central design choice is not to tune an intermediate query size, but to select an aggregation rule that determines whether the canonical algorithm is a sequential single-query scheme or a projection-based multi-query method.
Appendix
Appendix A High-Dimensional Behaviour of ZO-Align
In this appendix we analyze the behaviour of ZO-Align in the regime where the dimension $d$ is very large while the query batch size $q$ remains much smaller than $d$. Throughout, we retain the notation from the main text: $U\in\mathbb{R}^{d\times q}$ collects the $q$ query directions as columns, each sampled i.i.d. from $\mathcal{N}(0,I_{d})$, and
$P_{U}\;=\;U(U^{\top}U)^{-1}U^{\top}$
denotes the orthogonal projection onto the column space of $U$. In the idealized limit $\mu\to 0$, ZO-Align can be written as
$\hat{g}_{ALG}(x)\;=\;P_{U}\nabla f(x),$
so understanding the high-dimensional behaviour of ZO-Align reduces to understanding the random projector $P_{U}$.
A.1 Concentration of the Gram Matrix
It is convenient to view $U$ row-wise. Let $r_{1},\dots,r_{d}\in\mathbb{R}^{q}$ denote the rows of $U$. Since the columns of $U$ are i.i.d. $\mathcal{N}(0,I_{d})$, each row $r_{k}$ is an independent sample from $\mathcal{N}(0,I_{q})$ and
$U^{\top}U=\sum_{k=1}^{d}r_{k}^{\top}r_{k},\qquad\frac{1}{d}U^{\top}U=\frac{1}{d}\sum_{k=1}^{d}r_{k}^{\top}r_{k}$
is the empirical covariance matrix of $d$ i.i.d. Gaussian samples in $\mathbb{R}^{q}$.
The following standard result quantifies the concentration of this empirical covariance around the identity.
Lemma A.1.
There exist absolute constants $c,C>0$ such that the following holds. Let $U\in\mathbb{R}^{d\times q}$ have i.i.d. $\mathcal{N}(0,1)$ entries. Then for all $t\in(0,1)$,
$\Pr\!\left(\left\|\frac{1}{d}U^{\top}U-I_{q}\right\|_{op}\;\geq\;C\Big(\sqrt{\frac{q}{d}}+t\Big)\right)\;\leq\;2\exp(-cdt^{2}).$
In particular, if $d\gg q$ then with probability at least $1-2\exp(-c^{\prime}q)$,
$\left\|\frac{1}{d}U^{\top}U-I_{q}\right\|_{op}=O\!\left(\sqrt{\frac{q}{d}}\right).$
Proof.
This is a standard consequence of sample covariance concentration for isotropic sub-Gaussian vectors; see, e.g., the general treatments in Chapter 5 of (vershynin2018high). For completeness, we specialize the argument to the Gaussian case.
Row representation and reduction to one-dimensional marginals. Write $U$ row-wise as
$U=\begin{bmatrix}r_{1}^{\top}\\ \vdots\\ r_{d}^{\top}\end{bmatrix},\qquad r_{k}\in\mathbb{R}^{q},\quad r_{k}\sim\mathcal{N}(0,I_{q})\ \text{i.i.d.}$
Then
$\frac{1}{d}U^{\top}U=\frac{1}{d}\sum_{k=1}^{d}r_{k}r_{k}^{\top}$
is the empirical covariance matrix of $d$ i.i.d. Gaussian samples in $\mathbb{R}^{q}$.
Let
$\Sigma_{d}:=\frac{1}{d}U^{\top}U,\qquad A:=\Sigma_{d}-I_{q}.$
We need to control $\|A\|_{op}$. For any $x\in\mathbb{R}^{q}$ with $\|x\|=1$, we have
$x^{\top}Ax=x^{\top}\Sigma_{d}x-x^{\top}x=\frac{1}{d}\sum_{k=1}^{d}\langle r_{k},x\rangle^{2}-1.$
Define
$Z_{k}(x):=\langle r_{k},x\rangle,\qquad Y_{k}(x):=Z_{k}(x)^{2}-1.$
For fixed $x$ with $\|x\|=1$, $Z_{k}(x)\sim\mathcal{N}(0,1)$ i.i.d., so
$\sum_{k=1}^{d}Z_{k}(x)^{2}\sim\chi^{2}_{d}.$
Consequently, $\frac{1}{d}\sum_{k=1}^{d}Z_{k}(x)^{2}$ concentrates around $1$, and we can use a sharp tail inequality for the chi-square distribution due to Laurent and Massart (laurent2000adaptive).
Let $S(x):=\sum_{k=1}^{d}Z_{k}(x)^{2}$. By definition,
$x^{\top}Ax=\frac{1}{d}S(x)-1.$
Eq. 4.3-4.4 in (laurent2000adaptive) show that if $S\sim\chi^{2}_{d}$, then for all $u>0$,
$\Pr\!\big(S-d\geq 2\sqrt{du}+2u\big)$ $\leq e^{-u},$
$\Pr\!\big(d-S\geq 2\sqrt{du}\big)$ $\leq e^{-u}.$
Translating these inequalities to $\frac{1}{d}S(x)-1$ yields: for all $u>0$,
$\Pr\!\left(\frac{1}{d}S(x)-1\geq 2\sqrt{\frac{u}{d}}+2\frac{u}{d}\right)$ $\leq e^{-u},$
$\Pr\!\left(1-\frac{1}{d}S(x)\geq 2\sqrt{\frac{u}{d}}\right)$ $\leq e^{-u}.$
Taking $u\in(0,d)$, we have $\frac{u}{d}\leq 1$, so there exists an absolute constant $C_{0}>2$ such that
$2\sqrt{\frac{u}{d}}+2\frac{u}{d}\leq C_{0}\left(\sqrt{\frac{u}{d}}\right).$
Hence, for all $u\in(0,d)$,
$\Pr\!\left(\left|\frac{1}{d}S(x)-1\right|\geq C_{0}\sqrt{\frac{u}{d}}\right)\leq 2e^{-u}.$
Equivalently, for all $u\in(0,d)$,
$\Pr\!\left(\left|x^{\top}Ax\right|\geq C_{0}\sqrt{\frac{u}{d}}\right)\leq 2e^{-u}.$ (5)
Now restrict to thresholds $u=c_{1}dv^{2}$ with $v\in(0,1)$ and $c_{1}>0$ small enough so that $u<d$. Then
$\sqrt{\frac{u}{d}}=\sqrt{c_{1}}\,v,$
and (5) becomes
$\Pr\!\left(\left|x^{\top}Ax\right|\geq C_{1}v\right)\leq 2\exp(-c_{1}dv^{2}),$
for constants $C_{1}=C_{0}\sqrt{c_{1}}$ and $c_{1}>0$. Renaming $v$ to $u$ and $c_{1}$ to $c_{2}$, we obtain:
$\Pr\!\left(\left|x^{\top}Ax\right|\geq u\right)\leq 2\exp(-c_{2}du^{2}),\qquad\forall\,u\in(0,1),\ \forall\,x\in\mathbb{S}^{q-1},$ (6)
for some absolute constant $c_{2}>0$ and where $\mathbb{S}^{q-1}=\{x\in\mathbb{R}^{q}:\|x\|=1\}$.
Let $\mathbb{S}^{q-1}$ denote the unit sphere in $\mathbb{R}^{q}$. For $\varepsilon\in(0,1)$, an $\varepsilon$-net $\mathcal{N}_{\varepsilon}\subset\mathbb{S}^{q-1}$ is a finite set such that for any $x\in\mathbb{S}^{q-1}$ there exists $y\in\mathcal{N}_{\varepsilon}$ with $\|x-y\|\leq\varepsilon$. It is well known that there exists an $\varepsilon$-net with cardinality bounded by
$|\mathcal{N}_{\varepsilon}|\leq\left(1+\frac{2}{\varepsilon}\right)^{q}.$
We fix $\varepsilon=\frac{1}{4}$ and let $\mathcal{N}$ be a $\frac{1}{4}$-net; then
$|\mathcal{N}|\leq 9^{q}.$
We also use a standard net approximation for the operator norm of a symmetric matrix.
Claim A.2.
Let $A$ be a symmetric $q\times q$ matrix and let $\mathcal{N}$ be a $\frac{1}{4}$-net of $\mathbb{S}^{q-1}$. Then
$\|A\|_{op}=\sup_{x\in\mathbb{S}^{q-1}}|x^{\top}Ax|\leq 2\sup_{y\in\mathcal{N}}|y^{\top}Ay|.$
Proof of Claim A.2.
Let $x^{\star}\in\mathbb{S}^{q-1}$ satisfy $|{x^{\star}}^{\top}Ax^{\star}|=\|A\|_{op}$. Choose $y\in\mathcal{N}$ with $\|x^{\star}-y\|\leq\tfrac{1}{4}$. Then
$x^{\star}=y+(x^{\star}-y),\qquad\|x^{\star}-y\|\leq\tfrac{1}{4}.$
Using symmetry of $A$,
$x^{\star\top}Ax^{\star}=y^{\top}Ay+2(x^{\star}-y)^{\top}Ay+(x^{\star}-y)^{\top}A(x^{\star}-y).$
Taking absolute values and using $\|y\|,\|x^{\star}-y\|\leq 1$,
$\|A\|_{op}$ $\leq|y^{\top}Ay|+2\|x^{\star}-y\|\|A\|_{op}\|y\|+\|x^{\star}-y\|^{2}\|A\|_{op}$
$\leq|y^{\top}Ay|+\left(2\cdot\tfrac{1}{4}\cdot 1+\tfrac{1}{4}^{2}\right)\|A\|_{op}=|y^{\top}Ay|+\tfrac{9}{16}\|A\|_{op}.$
Rearranging gives
$\frac{7}{16}\|A\|_{op}\leq|y^{\top}Ay|,$
hence
$\|A\|_{op}\leq\frac{16}{7}\sup_{y\in\mathcal{N}}|y^{\top}Ay|\leq 2\sup_{y\in\mathcal{N}}|y^{\top}Ay|.$
Applying Claim A.2 to $A=\Sigma_{d}-I_{q}$ yields
$\left\|\frac{1}{d}U^{\top}U-I_{q}\right\|_{op}=\|A\|_{op}\leq 2\max_{y\in\mathcal{N}}|y^{\top}Ay|.$
Fix $\delta\in(0,1)$, to be chosen later. Using the net bound and (6), we have
$\Pr\!\left(\|A\|_{op}\geq\delta\right)$ $\leq\Pr\!\left(\exists\,y\in\mathcal{N}\text{ such that }|y^{\top}Ay|\geq\frac{\delta}{2}\right)$
$\leq\sum_{y\in\mathcal{N}}\Pr\!\left(|y^{\top}Ay|\geq\frac{\delta}{2}\right)$
$\leq|\mathcal{N}|\cdot 2\exp\!\left(-c_{2}d\frac{\delta^{2}}{4}\right)$
$\leq 2\exp\!\left(q\log 9-c_{3}d\delta^{2}\right),$
for some absolute constant $c_{3}>0$.
Now take
$\delta=C\Big(\sqrt{\frac{q}{d}}+t\Big),\qquad t\in(0,1),$
for a sufficiently large absolute constant $C>0$ (to be fixed). Then
$\delta^{2}=C^{2}\left(\frac{q}{d}+t^{2}+2t\sqrt{\frac{q}{d}}\right)\geq\frac{C^{2}}{2}\left(\frac{q}{d}+t^{2}\right),$
and hence
$q\log 9-c_{3}d\delta^{2}$ $\leq q\log 9-c_{3}d\cdot\frac{C^{2}}{2}\left(\frac{q}{d}+t^{2}\right)$
$=q\left(\log 9-\frac{c_{3}C^{2}}{2}\right)-\frac{c_{3}C^{2}}{2}dt^{2}.$
Choose $C$ large enough so that $\log 9-\frac{c_{3}C^{2}}{2}\leq-1$. Then
$q\left(\log 9-\frac{c_{3}C^{2}}{2}\right)\leq-q\leq 0,$
and hence
$q\log 9-c_{3}d\delta^{2}\leq-c_{4}dt^{2}$
for some absolute constant $c_{4}>0$. Plugging this into the previous bound gives
$\Pr\!\left(\left\|\frac{1}{d}U^{\top}U-I_{q}\right\|_{op}\geq C\Big(\sqrt{\frac{q}{d}}+t\Big)\right)\leq 2\exp(-c_{4}dt^{2}),$
which proves the first statement (with $c:=c_{4}$).
The in particular statement. For the second claim, take $t=\sqrt{q/d}$ (note that $t\in(0,1)$ as soon as $d$ is sufficiently large compared to $q$). Then
$C\Big(\sqrt{\frac{q}{d}}+t\Big)=2C\sqrt{\frac{q}{d}},$
and
$\Pr\!\left(\left\|\frac{1}{d}U^{\top}U-I_{q}\right\|_{op}\geq 2C\sqrt{\frac{q}{d}}\right)\leq 2\exp\!\left(-c_{4}d\frac{q}{d}\right)=2\exp(-c_{4}q).$
Thus, with probability at least $1-2\exp(-c_{4}q)$ we have
$\left\|\frac{1}{d}U^{\top}U-I_{q}\right\|_{op}\leq 2C\sqrt{\frac{q}{d}},$
which is the claimed $O(\sqrt{q/d})$ bound (with $c^{\prime}:=c_{4}$). ∎
Thus, in the regime $d\gg q$ the Gram matrix satisfies
$U^{\top}U=d\,(I_{q}+\Delta_{d}),\qquad\|\Delta_{d}\|_{op}=O\!\left(\sqrt{\frac{q}{d}}\right)$
with overwhelmingly high probability. Geometrically, the columns of $U$ are almost mutually orthogonal and have almost equal norms.
A.2 Approximate Structure of the Projection Operator
We now use Lemma A.1 to approximate the projection operator $P_{U}$ in the high-dimensional regime. Recall that
$P_{U}=U(U^{\top}U)^{-1}U^{\top}=U\big(d(I_{q}+\Delta_{d})\big)^{-1}U^{\top}=\frac{1}{d}U(I_{q}+\Delta_{d})^{-1}U^{\top}.$
Lemma A.3 (Approximation of $(U^{\top}U)^{-1}$).
Suppose $\|\Delta_{d}\|_{op}\leq\frac{1}{2}$. Then
$(I_{q}+\Delta_{d})^{-1}=I_{q}+E_{d},\qquad\|E_{d}\|_{op}\leq 2\|\Delta_{d}\|_{op},$
and hence
$(U^{\top}U)^{-1}=\frac{1}{d}(I_{q}+E_{d}),\qquad\|(U^{\top}U)^{-1}-\tfrac{1}{d}I_{q}\|_{op}\leq\frac{2}{d}\|\Delta_{d}\|_{op}.$
Proof.
If $\|\Delta_{d}\|_{op}\leq\frac{1}{2}$ then the Neumann series
$(I_{q}+\Delta_{d})^{-1}=\sum_{k=0}^{\infty}(-\Delta_{d})^{k}$
converges in operator norm, and
$(I_{q}+\Delta_{d})^{-1}-I_{q}=\sum_{k=1}^{\infty}(-\Delta_{d})^{k}.$
Using the triangle inequality and submultiplicativity yields
$\|E_{d}\|_{op}=\left\|\sum_{k=1}^{\infty}(-\Delta_{d})^{k}\right\|_{op}\leq\sum_{k=1}^{\infty}\|\Delta_{d}\|_{op}^{k}=\frac{\|\Delta_{d}\|_{op}}{1-\|\Delta_{d}\|_{op}}\leq 2\|\Delta_{d}\|_{op},$
where the last inequality uses $\|\Delta_{d}\|_{op}\leq\frac{1}{2}$. The bound for $(U^{\top}U)^{-1}$ follows by multiplication by $1/d$. ∎
Combining Lemmas A.1 and A.3, we obtain that for $d$ sufficiently large and $q\ll d$, with high probability,
[…]
We next propagate this approximation through to the projector $P_{U}$.
Proposition A.4 (Approximation of $P_{U}$ by $UU^{\top}/d$).
Let $U\in\mathbb{R}^{d\times q}$ have i.i.d. $\mathcal{N}(0,1)$ entries. There exist absolute constants $c,C>0$ such that the following holds. With probability at least $1-2\exp(-cq)$,
$\left\|P_{U}-\frac{1}{d}UU^{\top}\right\|_{op}\;\leq\;C\sqrt{\frac{q}{d}}.$
Consequently, for any $g\in\mathbb{R}^{d}$,
$\left\|P_{U}g-\frac{1}{d}UU^{\top}g\right\|\;\leq\;C\sqrt{\frac{q}{d}}\,\|g\|.$
Proof.
Write
$P_{U}-\frac{1}{d}UU^{\top}=U\Big((U^{\top}U)^{-1}-\tfrac{1}{d}I_{q}\Big)U^{\top}=UE_{d}U^{\top},$
where $E_{d}$ is as in Lemma A.3. Thus
$\left\|P_{U}-\frac{1}{d}UU^{\top}\right\|_{op}\leq\|U\|_{op}^{2}\,\|E_{d}\|_{op}.$
For a Gaussian matrix $U\in\mathbb{R}^{d\times q}$, standard random matrix bounds give
$\|U\|_{op}\leq\sqrt{d}+\sqrt{q}+t$
with probability at least $1-2\exp(-c^{\prime}t^{2})$ for some absolute constant $c^{\prime}>0$. Taking $t$ of order $\sqrt{q}$ and using $q\leq d$ yields $\|U\|_{op}^{2}=O(d)$ with probability at least $1-2\exp(-c^{\prime\prime}q)$ for some $c^{\prime\prime}>0$.
On the same high-probability event, Lemma A.1 ensures that $\|\Delta_{d}\|_{op}=O(\sqrt{q/d})$ and in particular $\|\Delta_{d}\|_{op}\leq\tfrac{1}{2}$ for $d$ sufficiently large. Lemma A.3 then gives
$\|E_{d}\|_{op}\leq 2\|\Delta_{d}\|_{op}=O\!\left(\sqrt{\frac{q}{d}}\right).$
Putting these bounds together,
$\left\|P_{U}-\frac{1}{d}UU^{\top}\right\|_{op}\leq C_{1}d\cdot C_{2}\sqrt{\frac{q}{d}}=C\sqrt{\frac{q}{d}}$
for suitable absolute constants $C_{1},C_{2},C>0$. The bound for $P_{U}g-(1/d)UU^{\top}g$ follows immediately by multiplying by $\|g\|$. ∎
Proposition A.4 formalizes the statement that, when $d\gg q$, the action of $P_{U}$ on any vector $g$ is very close (in norm) to the action of the simpler operator $(1/d)UU^{\top}$. In other words, within the sampled subspace, the update direction produced by ZO-Align is, up to a small relative error of order $\sqrt{q/d}$, the same as ZO-Avg.
A.3 Implications for ZO-Align in High Dimensions
The main advantage of ZO-Align over ZO-AVG is that it explicitly accounts for correlations among the sampled directions through the factor $(U^{\top}U)^{-1}$, and thus adjusts the relative contributions of different directions in $U$ according to the spectrum of $U^{\top}U$. When $d$ is moderate and $q$ is a nontrivial fraction of $d$, the Gram matrix $U^{\top}U$ can deviate substantially from a scaled identity, and these spectral corrections can significantly improve how the sampled subspace represents the local geometry of $f$.
In the high-dimensional regime $d\gg q$, however, Lemma A.1 and Proposition A.4 show that, with high probability,
$U^{\top}U\approx dI_{q},\qquad P_{U}\approx\frac{1}{d}UU^{\top},$
with approximation error of order $\sqrt{q/d}$ in operator norm. In this case the sampled directions are already almost mutually orthogonal with nearly equal norms, and the spectral corrections encoded in $(U^{\top}U)^{-1}$ reduce, to leading order, to a global scaling by $1/d$. As a consequence, the additional structural advantage of ZO-Align over naive aggregation inside the sampled subspace becomes much less pronounced when $d$ is extremely large and $q\ll d$.
It is important to emphasize that this high-dimensional geometric simplification does not alter the exact MSE formulas in Proposition 3.1, which are valid for every finite $d$. The high-dimensional analysis merely explains why the geometric distinction inside the sampled subspace becomes weaker as $d/q\to\infty$.
A.4 Numerical Illustration of $U^{\top}U\approx dI_{q}$
To empirically illustrate the concentration of the Gram matrix, we consider Gaussian query matrices $U\in\mathbb{R}^{d\times q}$ with a fixed $q=10$ and varying dimension $d\in\{10,100,1000,10000\}$. For each value of $d$, we sample $U$ with i.i.d. $\mathcal{N}(0,1)$ entries and compute the normalized Gram matrix
$G_{d}\;=\;\frac{1}{d}U^{\top}U\in\mathbb{R}^{q\times q}.$
We then visualize the absolute values $|G_{d}|$ as heatmaps. When $d$ is small, $G_{d}$ exhibits noticeable deviations from the identity, with substantial off-diagonal entries. As $d$ increases, the diagonal entries of $G_{d}$ stabilize near $1$, while the off-diagonal entries shrink toward 0, and the heatmaps become visually indistinguishable from those of an identity matrix.
A typical realization is shown in Figure 6. The progressive disappearance of off-diagonal structure as $d$ grows is consistent with Lemma A.1 and supports the approximation $U^{\top}U\approx dI_{q}$ used in the high-dimensional discussion above.
Figure 6: Heatmaps of the absolute value of the normalized Gram matrix $G_{d}=(1/d)U^{\top}U$ for $q=10$ and varying dimensions $d\in\{10,100,1000,10000\}$. Each $U$ has i.i.d. $\mathcal{N}(0,1)$ entries. As $d$ increases, $G_{d}$ becomes visually closer to the identity matrix, with diagonal entries near 11 and off-diagonal entries shrinking toward 0.
Appendix B Proofs
B.1 Proof of Proposition LABEL:prop:unbiasedness_zo_avg
@@
B.2 Proof of Proposition LABEL:prop:biasedness_zo_align
@@
B.4 Proof of Proposition 4.1
Proof.
From the $L$-smoothness of $f$, we have the descent lemma:
$f(x_{t+1})\leq f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{L}{2}\|x_{t+1}-x_{t}\|^{2}.$
For ZO-AVG, Substituting the update rule $x_{t+1}-x_{t}=-\eta_{t}\hat{g}_{AVG}(x_{t})$ and taking the expectation conditional on $x_{t}$:
Using the properties $\mathbb{E}[\hat{g}_{AVG}(x_{t})]=\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{AVG}(x_{t})\|^{2}]=\frac{q_{t}+d+1}{q_{t}}\|\nabla f(x_{t})\|^{2}$, we get:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{L\eta_{t}^{2}(q_{t}+d+1)}{2q_{t}}\|\nabla f(x_{t})\|^{2}.$
It is easy to see that the optimal step size $\eta_{t}^{*}=\frac{q_{t}}{L(q_{t}+d+1)}$.
Similarly, for ZO-Align, using the properties $\mathbb{E}[\hat{g}_{ALG}(x_{t})]=\frac{q_{t}}{d}\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{ALG}(x_{t})\|^{2}]=\frac{q_{t}}{d}\|\nabla f(x_{t})\|^{2}$, we have:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\frac{q_{t}}{d}\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{L\eta^{2}_{t}q_{t}}{2d}\|\nabla f(x_{t})\|^{2}.$
It is easy to see that the optimal step size $\eta_{t}^{*}=1/L$. ∎
B.5 Proof of Theorem 4.2
@@
Using the properties $\mathbb{E}[\hat{g}_{AVG}(x_{t})]=\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{AVG}(x_{t})\|^{2}]=\frac{q_{t}+d+1}{q_{t}}\|\nabla f(x_{t})\|^{2}$, we get:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{L\eta_{t}^{2}(q_{t}+d+1)}{2q_{t}}\|\nabla f(x_{t})\|^{2}.$
It is easy to see that the optimal step size $\eta_{t}^{*}=\frac{q_{t}}{L(q_{t}+d+1)}$. Plugging in the optimal step size, we obtain the one-step progress:
@@
By $\gamma$-strong convexity, $\|\nabla f(x_{t})\|^{2}\geq 2\gamma(f(x_{t})-f(x_{*}))$. Let $\delta_{t}=E[f(x_{t})-f(x_{*})]$. Taking the full expectation and applying the convexity condition gives:
$\delta_{t+1}\leq\left(1-\frac{\gamma q_{t}}{L(q_{t}+d+1)}\right)\delta_{t}$
@@
$\delta_{T}\leq\left(1-\frac{\gamma q_{T-1}}{L(q_{T-1}+d+1)}\right)\delta_{T-1}$
$\leq\left(1-\frac{\gamma q_{T-1}}{L(q_{T-1}+d+1)}\right)\left(1-\frac{\gamma q_{T-2}}{L(q_{T-2}+d+1)}\right)\delta_{T-2}$
$\dots\leq\left(\prod_{t=0}^{T-1}\left(1-\frac{\gamma q_{t}}{L(q_{t}+d+1)}\right)\right)\delta_{0}$
@@
For the second part, we must find the sequence $\{q_{t}\}_{t=0}^{T-1}$ that minimizes the final error bound under the budget constraint. This is equivalent to minimizing the total contraction factor. Let $\rho_{t}=\frac{\gamma q_{t}}{L(q_{t}+d+1)}$. We solve the following optimization problem:
@@
$\max_{\{q_{t}\},T}\sum_{t=0}^{T-1}\frac{\gamma}{L}h(q_{t})\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
@@
As in B.5, we establish the one-step contraction factor. Using the properties $\mathbb{E}[\hat{g}_{ALG}(x_{t})]=\frac{q_{t}}{d}\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{ALG}(x_{t})\|^{2}]=\frac{q_{t}}{d}\|\nabla f(x_{t})\|^{2}$, we have:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\frac{q_{t}}{d}\eta_{t}\|\nabla f(x_{t})\|^{2}+\frac{L\eta^{2}_{t}q_{t}}{2d}\|\nabla f(x_{t})\|^{2}.$
It is easy to see that the optimal step size $\eta_{t}^{*}=1/L$. Plugging in the optimal step size, we have:
@@
Applying the $\gamma$-strong convexity condition and taking the full expectation yields:
$E[f(x_{t+1})-f(x_{*})]\leq\left(1-\frac{\gamma q_{t}}{Ld}\right)(E[f(x_{t})]-f(x_{*}))$
@@
$\min_{\{q_{t}\},T}\prod_{t=0}^{T-1}\left(1-\frac{\gamma q_{t}}{Ld}\right)\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
Let $c=\gamma/(Ld)$. The objective function $F(q_{0},\dots,q_{T-1})=\prod_{t=0}^{T-1}(1-cq_{t})$ is a product of linear terms. This function is concave over the feasible region (the simplex defined by the budget constraint). A fundamental result in convex optimization is that the minimum of a concave function over a compact convex set is achieved at an extreme point (a vertex) of the set.